docia14:

Ronelle asked 100 students to choose their favorite subject from among mathematics, history, and art. She found that 22 students chose mathematics, 26 history, and the rest art. What percent of students chose art?

Gdub08:

26 52 students chose art. 52/100 = 13/25 13/25 = 0.52 0.52 = 52% 52%